OpenStudy (tumblewolf):

Write the slope-intercept form of the line that passes through the point (1, 0) and is parallel to x - y = 7. They didn't teach us what to when you have the parallel. Fan and medal to whoever figures it out!

OpenStudy (seascorpion1):

First you find the slope-intercept form of x-y=7, parallel lines have the same slope so your line will have the same slope as x-y=7. You are given a point so you should plug it into your equation y=mx+b (where m (the slope) is known) to work out b (the y-intercept) and then you will have found m and b so you've worked out the equation of the line in slope-intercept form. Does that help?